# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

### v1

@@ 1. Introduction
The prevalence of offensive language on social media platforms, such as Facebook and Twitter, has become a pressing concern in recent years. This surge can be attributed to the anonymity provided by these platforms and the lack of strict regulations and effective mechanisms to restrain such behaviour (Fortuna and Nunes, 2018). While social media platforms have fostered connections and bridged global distances, they have in turn enabled the spread of hate speech and various forms of offensive language (Chhabra and Vishwakarma, 2023). Offensive language broadly refers to expressions that may upset or annoy others, while hate speech escalates beyond mere offensive behaviour to inciting discrimination, hostility, or violence against individuals or groups based on their race, religion, ethnicity, gender, or other identity factors (for Peace, [n. d.]). Hate speech is considered a human rights violation in many legal systems due to its potential to cause real-world harm and perpetuate discrimination against targeted groups (UNSPAHS, 2019). Hence, there is a growing scholarly focus on analysing and developing approaches to mitigate hate speech in addition to offensive content.
The Natural Language Processing (NLP) community has proposed a variety of research techniques to detect offensive content based on both traditional machine learning and advanced neural network-based approaches. While early work on offensive language detection focused on monolingual settings (i.e. English), research in recent years has increasingly focused on tackling the problem in multilingual settings to broaden the ability to deal with more languages, a key requirement given the presence of diverse languages in social media platforms (Pamungkas et al., 2021b). This however generally hindered by the limited availability of labelled data and the high variability of offensive language across diverse cultures and languages (Plaza-del arco et al., 2023).
@@
The earliest work in our survey is from 2018 (Mathur et al., 2018b) and, since then, interest in cross-lingual offensive language detection has increased, but is yet to be covered in a comprehensive survey. Existing surveys concentrate on monolingual detection or specific characteristics of offensive language, while the unique CLTL application in offensive language detection is seldom considered. Relevant surveys exist (Pamungkas et al., 2021b; Chhabra and Vishwakarma, 2023), but their scope goes beyond cross-lingual scenarios, also covering cross-domain and multimodal dimensions, and hence lacking our survey’s focus on CLTL.
@@ 1.1. Scope of the Survey
Our objective is to systematically review existing literature, elucidating different CLTL techniques employed in the task of offensive language detection. Consequently, this survey discusses 67 papers in this area. This is the first systematic and holistic overview of recent studies, challenges and open research avenues of using CLTL to detect offensive language in social media across languages. We describe those studies according to multiple aspects, including the multilingual datasets employed, cross-lingual resources leveraged, levels of transfer, and cross-lingual strategies applied. Inspired by Pikuliak et al. (2021), we outline three CLTL transfer approaches based on the level at which knowledge transfer occurs across languages, namely instance transfer, feature transfer, and parameter transfer. We also illustrate prevalent transfer strategies utilised based on these transfer levels. In addition, we highlight challenges and posit potential opportunities for future research in this field. Furthermore, we make our survey resources available online,111https://github.com/aggiejiang/crosslingual-offensive-language-survey including two comprehensive tables that provide accessible references to the multilingual datasets and CLTL methods used in surveyed papers to facilitate easy comparison and discovery of related works.
@@ 1.2. Survey Structure
The rest of the paper is organised as follows. Section 2 describes the definitions of offensive language and the framework of cross-lingual detection, and distinguishes ‘cross-lingual’ from related terminologies such as ‘multilingual’. Related surveys are also summarised. Section 3 summarises our survey methodology. Then we summarise and analyse the multilingual datasets leveraged in surveyed papers in Section 4. In Section 5, we describe diverse linguistic resources and tools used in cross-lingual studies. Section 6 discusses the three transfer levels of CLTL in offensive language detection, complemented by transfer strategies based on these layers. Section 7 presents current challenges and future research directions in this area, concluding the survey in Section 8.
@@ 2.1. Cross-linguality in offensive Language Detection
2.1.1. Task Definition and Workflow
The task of cross-lingual hate speech detection focuses on the identification and categorisation of offensive or hateful text across different languages. The primary objective for CLTL is to leverage the knowledge acquired from one language (source language with more resources) to enhance the hate speech detection in another language (target language with less resources), especially when the labelled data in the target language is scarce or unavailable (Pamungkas et al., 2021b). The workflow of cross-lingual hate speech is presented below:
Data Preparation. Let and represent the source and target languages. Then let and represent the source and target datasets respectively, where denotes text data and denotes labels indicating that the text is offensive or not. The cross-lingual task contains a predictive function , a feature space , and a label space , where 0 denotes non-offensive and 1 denotes offensive content.
Cross-lingual Training. The training stage for cross-lingual transfer can use only or both and . For only as training data, a model is trained to learn a function , where is the feature space of . Then adapt the learned function to a function , where is the feature space of , with minimal loss in performance. For both and as training data, a model is trained to learn a function , where is the joint feature space of . Various cross-lingual training strategies could be employed for cross-lingual transfer to obtain , where more details are described in Section 6.
Model Adaptation. If label is available for , fine-tune on full or few shots of to obtain a model , further adapting the function to the target domain: .
Detection on Target Language. Apply predictive functions or to to detect offensive content in and predict labels: or .
2.1.2. Cross-lingual & Multilingual & Code-mixing
While “cross-lingual learning” typically denotes the application of CLTL techniques, the term “multilingual learning” is sometimes used interchangeably (Pikuliak et al., 2021). Some multilingual data may contain code-mixed content, i.e. a mixture of languages in the same sentence (Ibrohim and Budi, 2019a). Although some researchers draw parallels between these terms and the relationship between transfer learning and multi-task learning (Wang and Zheng, 2015), we adopt a broader perspective on cross-lingual learning. We define cross-lingual learning as the process of transferring knowledge between different languages, a concept that inherently subsumes multilingual learning.
@@ 2.2. Related Surveys
General Offensive Language Detection. Surveys on offensive language detection provide general insights into the field (Schmidt and Wiegand, 2017; Fortuna and Nunes, 2018; Jahan and Oussalah, 2023; Mishra et al., 2019). Key works by Schmidt and Wiegand (2017) and Fortuna and Nunes (2018) stand out as the two most cited reviews in the NLP community, providing definitions of hate speech and related terms, summarising previous methods, and discussing the existing limitations and challenges. Some of these surveys focus on in-depth discussions of specific aspects when reviewing studies in hate speech detection, such as data sources (Alkomah and Ma, 2022; Vidgen and Derczynski, 2020; Poletto et al., 2021), methodologies (Alkomah and Ma, 2022; Ayo et al., 2020; Demilie and Salau, 2022), and challenges (Ayo et al., 2020; Kiritchenko et al., 2021; Yin and Zubiaga, 2021; Garg et al., 2023). Additionally, several surveys have also delved into related areas like cyberbullying (Yi and Zubiaga, 2023) and the intersection of hate speech and fake news (Demilie and Salau, 2022), and others review ethical considerations in the development of hate speech detection (Vidgen and Derczynski, 2020; Kiritchenko et al., 2021).
Cross-lingual and Multilingual Offensive Language Detection. Several surveys focus on studies in single but non-English languages, such as Arabic (Husain and Uzuner, 2021; Al-Hassan and Al-Dossari, 2019), Indian (Dowlagar and Mamidi, 2021) and Ethiopian (Demilie and Salau, 2022) languages. Several surveys for general hate speech detection also provide a brief summary of multilingual or cross-lingual contributions in this area (Poletto et al., 2021; Al-Hassan and Al-Dossari, 2019; Yin and Zubiaga, 2021). However, systematic reviews focusing explicitly on cross-lingual and multilingual perspectives remain scarce (Pamungkas et al., 2021b; Chhabra and Vishwakarma, 2023). Pamungkas et al. (2021b) provide an overview of available datasets and approaches in cross-domain and cross-lingual settings. Chhabra and Vishwakarma (2023) conduct a comprehensive review of multimodal and multilingual hate speech detection, including benchmark datasets, state-of-the-art methods, challenges, and the latest evaluation techniques.
Our survey instead focuses on cross-lingual and multilingual offensive language detection by providing a systematic review into available datasets, cross-lingual resources, existing CLTL approaches, related challenges and future research directions.
@@ 3. Survey Methodology
3.1. Selection Process
Search Databases
We retrieve publications from scientific databases including ACL Anthology,222https://aclanthology.org/ Google Scholar,333https://scholar.google.com/ and DBLP Computer Science Bibliography.444https://dblp.org/
Keyword Selection
We create a list of 99 pairwise combinations of words in (i) and (ii): (i) 9 keywords related to offensive language, such as “hate”, “abus*”, “offens*”, “toxic”, “cyberbully”, “gender”, “sexis*”, “misogyn*”, “rac*”; (ii) 11 keywords related to multilingual and cross-lingual scenarios, such as “cross-lingual”, “multilingual”, “low resource”, “under resource”, “cross language”, “between language”, “cross linguistic”, “transfer”, “code switch”, “zero shot”, “few shot”.
Search Process
In addition to studies matching the 99 keyword pairs above, we also identify and include 5 additional studies cited in relevant surveys. This search was conducted on July 29th, 2023.
Eligibility Criteria
Following removal of duplicates and irrelevant papers, the remainder 455 publications were checked against the inclusion / exclusion criteria in Table 1.
@@ 3.1. Selection Process
(i) Describe cross-lingual approaches to detect general abusive content (toxic, hate, etc.) |
(i) Study transfer across dimensions other than languages (e.g. platforms, domains, datasets) |
(ii) Describe cross-lingual approaches to detect subtypes of abusive content (e.g. sexism, racism) |
(ii) Study low-resource languages without knowledge transfer approaches |
(iii) Describe multilingual approach with knowledge transfer between languages |
(iii) Relevant studies from competition reports and papers |
(iv) Study text classification cases in cross-lingual hate speech scenarios |
(iv) Master or PhD dissertations |
3.2. Summary of Surveyed Studies
In the end we obtained 67 relevant papers, with their temporal distribution shown in Figure 2 and showing a growing trend from the first publication in 2018 to 2022.
@@ 4. Multilingual Hate Speech Datasets
We investigate 82 datasets utilised in existing surveyed studies on cross-lingual hate speech detection across different aspects. We summarise these datasets and their characteristics in Table 2.

### v2

# Cross-lingual Offensive Language Detection: A Systematic Review of Dataset, Approach and Challenge

@@ 1. Introduction
The prevalence of offensive language on social media platforms, such as Facebook and Twitter, has become a pressing concern in recent years. This surge can be attributed to the anonymity provided by these platforms and the lack of strict regulations and effective mechanisms to restrain such behaviour (Fortuna and Nunes, 2018). While social platforms have fostered connections and bridged global distances, they have in turn enabled the spread of hate speech (HS) and various forms of offensive language (Chhabra and Vishwakarma, 2023). Offensive language broadly refers to expressions that may upset or annoy others, while HS escalates beyond mere offensive behaviour to inciting discrimination, hostility, or violence against individuals or groups based on their race, religion, ethnicity, gender, or other identity factors (for Peace, [n. d.]). HS is considered a human rights violation in many legal systems due to its potential to cause real-world harm and perpetuate discrimination against targeted groups (UNSPAHS, 2019). Hence, there is a growing scholarly focus on developing approaches to mitigate HS in addition to offensive content.
The Natural Language Processing (NLP) community has proposed a variety of research techniques to detect offensive content based on both traditional machine learning and advanced neural network-based approaches. While early work focused on monolingual settings (i.e. English), research in recent years has increasingly focused on tackling the problem in multilingual settings to broaden the ability to deal with more languages, a key requirement given the presence of diverse languages in social platforms (Pamungkas et al., 2021b). This however generally hindered by the limited availability of labelled data and the high variability of offensive language across diverse cultures and languages (Plaza-del arco et al., 2023).
@@
Interest in cross-lingual offensive language detection has increased since the first work in 2018 (Mathur et al., 2018b), but this is yet to be covered in a comprehensive survey. Existing surveys concentrate on monolingual detection or specific characteristics of offensive language, while the unique CLTL application in offensive language detection is seldom considered. Relevant surveys exist (Pamungkas et al., 2021b; Chhabra and Vishwakarma, 2023), but their scope goes beyond cross-lingual scenarios, also covering cross-domain and multimodal dimensions, and hence lacking our survey’s focus on CLTL.
@@ 1.1. Scope of the Survey
Our objective is to provide a systematic review of existing literature on CLTL techniques for offensive language detection, and consequently, this survey discusses 67 papers in this area. To make this scope explicit, we set out the following key objectives:
Map the landscape of multilingual datasets and cross-lingual resources used in offensive language detection, discussing aspects such as languages, label schemes, and availability.
Classify and analyse methods by introducing a fine-grained taxonomy of CLTL approaches tailored to this task, distinguishing instance-, feature-, and parameter-level transfer.
Identify key challenges that limit progress, with a focus on language diversity, dataset scarcity, cultural variation, and modelling limitations in low-resource scenarios.
Recommend future research directions for developing more generalisable, robust, and ethically sound cross-lingual detection systems.
Furthermore, we make survey resources available online,111https://github.com/aggiejiang/crosslingual-offensive-language-survey including two comprehensive tables that summarise multilingual datasets and CLTL methods used in surveyed papers.
@@ 1.2. Survey Structure
The rest of the paper is organised as follows. §2 describes the definitions of offensive language and the framework of cross-lingual detection, and distinguishes ‘cross-lingual’ from related terminologies such as ‘multilingual’. Related surveys are also summarised. §3 summarises our survey methodology. Then we summarise and analyse the multilingual datasets leveraged in surveyed papers in §4. In §5, we describe diverse linguistic resources and tools used in cross-lingual studies. §6 discusses the three transfer levels of CLTL in offensive language detection, complemented by transfer strategies based on these layers. §7 presents current challenges and future research directions in this area, concluding the survey in §8.
@@ 2.1. Cross-linguality in Offensive Language Detection
Task Definition and Workflow. The task focuses on the identification and categorisation of offensive or hateful text across different languages. The primary objective for CLTL is to leverage the knowledge acquired from one language (source language with more resources) to enhance the HS detection in another language (target language with less resources), especially when the labelled data in the target language is scarce or unavailable (Pamungkas et al., 2021b). The task workflow can be thought of as a four-stage process:
Data Preparation. Gather text in both a source language with more labelled examples and a target language with fewer or no labels. The source data provides the foundation for learning, while the target data is where we ultimately want to detect offensive content.
Cross-lingual Training. Use the source data alone, or combine it with any available target data, to train a model how to recognise offensive language patterns.
Model Adaptation. If any labelled examples exist in the target language, fine-tune the model so it better reflects specific characteristics of the target language.
Detection on Target Language. Apply the adapted model to unseen target language text to assign offensive/non-offensive labels.
Formally, we represent the process as follows. Let and be the source and target languages, with datasets and , where denotes text and denotes non-offensive (0) or offensive (1) labels. The cross-lingual task seeks a function mapping a feature space to the target label space. In the training stage, it can use only or both and :
Source-only training: Train on to learn , then adapt to .
Joint training: Train on to learn based on joint feature space .
If is available, fine-tune (or ) on full or few-shot , producing for prediction:
Cross-lingual & Multilingual & Code-mixing. While “cross-lingual learning” typically denotes the application of CLTL techniques, the term “multilingual learning” is sometimes used interchangeably (Pikuliak et al., 2021). Some multilingual data may contain code-mixed content, i.e. a mixture of languages in the same sentence (Ibrohim and Budi, 2019a). Although some researchers draw parallels between these terms and the relationship between transfer learning and multi-task learning (Wang and Zheng, 2015), we adopt a broader perspective on cross-lingual learning. We define cross-lingual learning as the process of transferring knowledge between different languages, a concept that inherently subsumes multilingual learning.
@@ 2.2. Related Surveys
General Offensive Language Detection. Surveys on offensive language detection provide general insights into the field (Schmidt and Wiegand, 2017; Fortuna and Nunes, 2018; Jahan and Oussalah, 2023; Mishra et al., 2019). Key works by Schmidt and Wiegand (2017) and Fortuna and Nunes (2018) stand out as the two most cited reviews in the NLP community, providing definitions of HS and related terms, summarising previous methods, and discussing the existing limitations and challenges. Some of these surveys focus on in-depth discussions of specific aspects when reviewing studies in HS detection, such as data sources (Alkomah and Ma, 2022; Vidgen and Derczynski, 2020; Poletto et al., 2021), methodologies (Alkomah and Ma, 2022; Ayo et al., 2020; Demilie and Salau, 2022), and challenges (Ayo et al., 2020; Kiritchenko et al., 2021; Yin and Zubiaga, 2021; Garg et al., 2023). Additionally, several surveys have also delved into related areas like cyberbullying (Yi and Zubiaga, 2023) and the intersection of HS and fake news (Demilie and Salau, 2022), and others review ethical considerations in the development of HS detection (Vidgen and Derczynski, 2020; Kiritchenko et al., 2021).
Cross-lingual and Multilingual Offensive Language Detection. Several surveys focus on studies in single but non-English languages, such as Arabic (Husain and Uzuner, 2021; Al-Hassan and Al-Dossari, 2019), Indian (Dowlagar and Mamidi, 2021) and Ethiopian (Demilie and Salau, 2022) languages. Surveys for HS detection also provide a brief summary of multilingual or cross-lingual contributions (Poletto et al., 2021; Al-Hassan and Al-Dossari, 2019; Yin and Zubiaga, 2021). However, systematic reviews focusing explicitly on cross-lingual perspectives remain scarce (Pamungkas et al., 2021b; Chhabra and Vishwakarma, 2023). Pamungkas et al. (2021b) provide an overview of datasets and approaches in cross-domain and cross-lingual settings. Chhabra and Vishwakarma (2023) conduct a comprehensive review of multimodal and multilingual HS detection, including benchmark datasets, state-of-the-art methods, challenges, and evaluation techniques.
Strength of our survey. In contrast, our work presents:
Fine-grained taxonomy of CLTL approaches for offensive language detection, structured into instance-, feature-, and parameter-level transfer, each with detailed strategy categorisation;
Comprehensive dataset library including rich metadata (e.g., language families, annotation schemes, abuse types, availability, etc.);
In-depth analysis of challenges unique to cross-lingual offensive language detection from aspects of languages, data sources and approaches;
Future-oriented perspective that synthesises promising research directions and provides recommendations for future research.
@@ 3. Survey Methodology
Search Databases. We retrieve publications from scientific databases including ACL Anthology,222https://aclanthology.org/ Google Scholar,333https://scholar.google.com/ and DBLP Computer Science Bibliography.444https://dblp.org/
Keyword Selection. We create a list of 99 pairwise combinations of words in (i) and (ii): (i) 9 keywords related to offensive language, such as “hate”, “abus*”, “offens*”, “toxic”, “cyberbully”, “gender”, “sexis*”, “misogyn*”, “rac*”; (ii) 11 keywords related to multilingual and cross-lingual scenarios, such as “cross-lingual”, “multilingual”, “low resource”, “under resource”, “cross language”, “between language”, “cross linguistic”, “transfer”, “code switch”, “zero shot”, “few shot”.
Search Process. In addition to studies matching the 99 keyword pairs above, we also identify and include 5 additional studies cited in relevant surveys. This search was conducted on July 29th, 2023.
Eligibility Criteria. Following removal of duplicates and irrelevant papers, the remainder 455 publications were checked against the inclusion / exclusion criteria in Table 1.
@@
| (i) Describe cross-lingual approaches to detect general abusive content (toxic, hate, etc.) | (i) Study transfer across dimensions other than languages (e.g. platforms, domains, datasets) |
| (ii) Describe cross-lingual approaches to detect subtypes of abusive content (e.g. sexism, racism) | (ii) Study low-resource languages without knowledge transfer approaches |
| (iii) Describe multilingual approach with knowledge transfer between languages | (iii) Relevant studies from competition reports and papers |
| (iv) Study text classification cases in cross-lingual HS scenarios | (iv) Master or PhD dissertations |
Summary. In the end we obtained 67 relevant papers, with their temporal distribution shown in Figure 2 and showing a growing trend from the first publication in 2018 to 2022.
@@ 4. Multilingual Hate Speech Datasets
We investigate 82 datasets utilised in existing surveyed studies on cross-lingual HS detection across different aspects. We summarise these datasets and their characteristics in Table LABEL:tab:datasets.